The molar heat capacity of a gas is given in the table below: T(K) 373 29.19 473 29.29 573 29.46 673 29.68 773 29.97 873 30.27 973 30.56 c_{D} (J/mol.K) a) Fit a 2 ^ (nd) order polynomial to the data to find c_{p} as a function of Temperature. Use the Cramer's Rule to solve the equations. Show all steps. b) Find the value of the heat capacity at 625 K using the equation derived in part a.


1
Expert's answer
2022-01-10T14:36:03-0500


1) "\Delta E=nCvdT"

n=3

T2=50oC=323,15K

T1=0oC=273.15K

"\Delta E=nCvdT=3\times5\times(323.15-273.15)=750 cal"


2)

"\Delta H=nCpdT=n(Cv+R)dT"

"\Delta H=n(Cv+R)dT=3\times(5+2)\times(323.15-273.15)=1050 cal"